# Learning to Edit Visual Programs with Self-Supervision

### v1

@@ Appendix D Editing Operations
We once again note that each of these edit operations has a local effect. For instance, as depicted in Figure 1 adding a new transform function inserts a transform node into an already existing tree of functions. Similarly, removing a transform functions simply results in forming a skip connection from the chosen operator’s parent function to the chosen operator’s child function. Somewhat more arbitrary changes can be enacted by removing or adding combinators, in order to produce or remove entire expression trees, though these are inserted or removed from specific locations.
D.1 findEdits Algorithm
@@
As we mention in the discussion, while it may be possible to employ a more general algorithm in this step, all of our experiments use domain-aware reasoning to parse out a set of edit operations from edit types we detail above. Below we provide the high-level intuition behind our procedure.
@@
From the above procedure we know the edit operations and the edit cost of transforming any sub-expression $a$ into another sub-expression $b$. We design our DSLs so that it is possible to break each program into a series of sub-expressions. For Layout this is done by splitting the top-level UBlocks into the top-level ShBlocks. For CSG this is done by splitting each POS block into E blocks and each NEG block int E blocks. Note that not there is some order dependency in this match: for CSG positive sub-expressions must be matched to other positive sub-expressions, while negative sub-expressions must be matched to other negative sub-expressions. For the Layout domain, Union is not an order invariant operator as it controls how primitives are layered on the canvas. Therefore we keep the order of Layout sub-expressions fixed, although we allow each sub-expression to optionally match to an empty sub-expression $\emptyset$. A match from $a_{i}$ to $\emptyset$ implies that $a_{i}$ will be removed with a RC edit operation, while a match from $\emptyset$ to $b_{i}$ implies that $b_{i}$ will be added with a AC edit operation. We consider all valid possible ways to enact this matching by calculating the cost of each sub-expression match and then extracting out a solution with the Hunagarian matching algorithm [22].
D.2 Converting edits operations to training data
@@
Appendix E Program Corruption
@@
While this variant is not as a performant as our default version, it still sources useful training data for our edit network. Our view is that, when possible, it is better to source these edit operations by considering start program and end program pairs, but for domains where such edit differences scripts are hard analytically find, this corruption variant offers an alternative. While its possible that better corruption processes could close this gap, designing them is non-trivial. From one perspective, when we want to combine one-shot models and edit networks at inference time, the corruption behavior we want should noise ‘end’ programs towards those produced by the one-shot model – this is exactly the distribution we get access to with the findEdits approach that considers program-to-program transformations. Another benefit of this formulation, is that the distribution of edit operations we train over is naturally allowed to evolve and keeps in sync automatically with the finetuned one-shot model. Maintaining this property with a corruption-based procedure would likely be impractical.

### v2

@@ E.1 Local Edit Operations
  • Modify Combinator (MC): modifies a combinator function (e.g. changing difference to an intersection). Requires additional parameter predictions to set the new function.
@@
We once again note that each of these edit operations has a local effect. For instance, as depicted in Figure 1 adding a new transform function inserts a transform node into an already existing tree of functions. Similarly, removing a transform functions simply results in forming a skip connection from the chosen operator’s parent function to the chosen operator’s child function. Somewhat more arbitrary changes can be enacted by removing or adding combinators, in order to produce or remove entire expression trees, though these are inserted or removed from specific locations. While this framing does focus on local edits, and as such our edit network makes local changes in program space, some of these changes can have dramatic effects in the execution space. For instance, consider changing a boolean operation type in CSG from difference to union.
E.2 findEdits Algorithm
@@
We design our findEdits algorithm to try to find the “minimal cost” set of edit operations that would transform a start program to an end program. Our instantiation of the algorithm works over multiple visual programming domains for the set of edit operations we consider. However, there are many alternative ways this algorithm could be instantiated, and such alterations could prove useful in helping our method adapt for very different domains. As one extreme point, consider that for general purpose programming languages, a simple “git-diff” command could be used to turn a (start, end) program pair into a set of local insert/keep/delete edits.
Our implementation evaluates valid transformations in terms of program semantics (e.g. executions) not just syntax (e.g. token matching), as there are many distinct programs in our domains that will produce equivalent execution outputs (e.g. note that the argument ordering of union for CSG languages does not change the execution output). We hypothesize that using a findEdit algorithm alternative that does not consider such "semantic-equivalences" would result in a “worse” edit network (as the patterns in the training data would be less consistent), but it would be interesting to explore how different algorithms would effect system performance in future work.
@@
From the above procedure we know the edit operations and the edit cost of transforming any sub-expression $a$ into another sub-expression $b$. We design our DSLs so that it is possible to break each program into a series of sub-expressions. For Layout this is done by splitting the top-level UBlock into the top-level ShBlocks. For CSG this is done by splitting each POS block into E blocks and each NEG block into E blocks. Note that there is some order dependency in this match: for CSG positive sub-expressions must be matched to other positive sub-expressions, while negative sub-expressions must be matched to other negative sub-expressions. For the Layout domain, Union is not an order invariant operator as it controls how primitives are layered on the canvas. Therefore we keep the order of Layout sub-expressions fixed, although we allow each sub-expression to optionally match to an empty sub-expression $\emptyset$. A match from $a_{i}$ to $\emptyset$ implies that $a_{i}$ will be removed with a RC edit operation, while a match from $\emptyset$ to $b_{i}$ implies that $b_{i}$ will be added with a AC edit operation. We consider all valid possible ways to enact this matching by calculating the cost of each sub-expression match and then extracting out a solution with the Hungarian matching algorithm [22].
E.3 Converting edits operations to training data
@@
E.4 Generality of our framing
While we designed our edit operations with the task of visual program induction in mind, we believe that these operations are quite flexible. Many other functional DSLs for visual programs (and for other program synthesis tasks) could likely be subsumed directly under our framework, as long as these languages meet the criteria described in Section E.1. For instance, this set of edit operations should be able to handle any DSL expressible as a Context Free Grammar.
Under these assumptions, the edit operations we use are quite basic and make limited domain assumptions. For an input functional program, edits to transform functions allow for local edits (delete/insert/modify) that don’t affect the branching factor, while edits to combinator functions allow for local edits (delete/insert) that do affect the branching factor. We employ this formulation for a few reasons: (1) it is general enough to support any program-program transformation (under our assumption set) and (2) applying any of these local edits creates a syntactically complete program that can be immediately executed.
That said, our framework and core contributions are not tied to this specific set of edit operations. Our edit network and proposed training scheme could be easily integrated with any set of local edit operations (assuming an analogous findEdits algorithm can be designed for this new set of edits). So while we believe that the set of edit operations we introduce is quite general (as evidenced by their usefulness across multiple difficult visual programming domains), we are also excited to see how our general learning-to-edit framework could be extended to even more complex DSLs and edit operations.
Appendix F Program Corruption
@@
While this variant is not as a performant as our default version, it still sources useful training data for our edit network. Our view is that, when possible, it is better to source these edit operations by considering start program and end program pairs, but for domains where such edit difference scripts are hard to analytically find, this corruption variant offers an alternative. While its possible that better corruption processes could close this gap, designing them is non-trivial. From one perspective, when we want to combine one-shot models and edit networks at inference time, the corruption behavior we want should noise ‘end’ programs towards those produced by the one-shot model – this is exactly the distribution we get access to with the findEdits approach that considers program-to-program transformations. Another benefit of this formulation, is that the distribution of edit operations we train over is naturally allowed to evolve and keeps in sync automatically with the finetuned one-shot model. Maintaining this property with a corruption-based procedure would likely be impractical.